# Around the codifferential of products of p-forms and generalized interior products

### v1

@@ I Introduction
Computations with $p$-forms involve differential operators such as the exterior derivatives $d$ and the Lie derivatives $\mathscr{L}_{v}$, both of which fulfilling a Leibniz identity with respect to the exterior product $\wedge$. In stark contrast is the codifferential $\delta$ which is built out of $d$ and the Hodge operator * and for which there is no such identity. However, some other identities involving $\delta$ can be found scattered here and there in the literature, often in a wider setting. This short paper aims to gather such formulas in one compact space making it a useful toolbox for the practitioners of the field. Those formulas are extended further using a generalized interior product, producing new (up to our knowledge) formulas such as Eqs. (60)–(64).
Section II deals with the relationship between the codifferential $\delta$ and the Schouten-Nijenhuis bracket brought back to forms, facts known for a long time once cast in the framework of Gertenhaber algebras. Some explicit and little known formulas are gleaned along the way. In Sec. III the interior product $i_{v}$ is generalized in a somewhat natural manner, it provides a set of new identities of their own and in peculiar it involves naturally the Schouten-Nijenhuis bracket. A set of noteworthy formulas involving the differential operators $d$, $\mathscr{L}_{v}$, $\delta$ and $\square$ are recorded. Our conventions regarding differential geometry follow those of Ref. 1 and needed formulas are briefly recalled in appendix A. The outline of some longer calculations are gathered in appendix B in order not to clutter the main body of the paper.
@@ II The Schouten-Nijenhuis bracket on forms and its properties
Schouten[2, 3], and then Schouten-Nijenhuis[4, 5], bracket came from the natural endeavor in differential geometry of generalizing the Lie derivative not just along vector but also along tensors. In particular it is the unique extension generalizing the Lie derivative on multivectors[6, 7] (i.e. the sections of $\bigwedge^{p}TM$). In addition to this, it fulfills a wealth of algebraic relations making it one of the simplest incarnation of a Gerstenhaber algebra[8, 9, 10]. In this section we expose how, once brought back to forms, it arises out of the codifferential of the product of two forms. The bracket can be viewed from many angles[11, 12], in particular it plays a key role in Poisson geometry[13, 14]. For additional material regarding Schouten brackets and its generalizations we refer the reader to Ref. 15 for a beautiful overview.
@@
$\llbracket\lambda,\alpha\rrbracket=\flat(\mathscr{L}_{\sharp\lambda}\sharp\alpha),$ $\lambda\in\Omega^{1},\alpha\in\Omega^{a}.$ (8)
@@
$\llbracket\lambda,\alpha\rrbracket=(\mathscr{L}_{\sharp\lambda}\tilde{g})^{ab}j_{a}i_{b}\alpha+\mathscr{L}_{\sharp\lambda}\alpha,$ (9)
where $\tilde{g}$ is the corresponding metrics on forms and $j_{a}$, $i_{b}$ stands for $j_{e_{a}}$ and $i_{e_{b}}$. When $\sharp\lambda$ is a Killing vector it simplifies to $\llbracket\lambda,\alpha\rrbracket=\mathscr{L}_{\sharp\alpha}\beta$.
@@
and makes the Gerstenhaber algebra an exact Gerstenhaber algebra, that is a Batalin-Vilkovisky algebra[18, 19]. Equation (29), or Eq. (19), has been known for a long time[20, 21] (see also Ref. 14) and can be recovered from Gerstenhaber original work[8] through Hochschild-Kostant-Rosenberg theorem[22], or be found by other means[23]. As the bracket is generated by $\delta$ it behaves nicely with respect to it, namely:
@@
$\llbracket\alpha,\beta\rrbracket=-(\nabla_{a}\alpha)\wedge i^{a}\beta+(-1)^{a+1}(i^{a}\alpha)\wedge\nabla_{a}\beta,$ (31)
@@ III A generalized interior product and its interplay with the Schouten-Nijenhuis bracket
for $\alpha\in\Omega^{a}$, $\beta\in\Omega^{b}$ and $\gamma\in\Omega^{c}$ forms, $\phi$ a function and $\omega$ the volume-form. For $\alpha,\beta\in\Omega^{a}$ two forms of the same degree one recovers the product between $p$-forms:
@@ Appendix A Overview of needed formulas of differential geometry
We recall that the interior product appears in the Lie derivative through Cartan formula $\mathscr{L}_{v}=i_{v}d+di_{v}$.
@@
see say App. A of Ref. 24 for additional identities. The generalized interior product stems from Eq. (75).

### v2

@@ I Introduction
Computations with $p$-forms involve differential operators such as the exterior derivatives $d$ and the Lie derivatives $\mathscr{L}_{v}$, both of which fulfilling a Leibniz identity with respect to the exterior product $\wedge$. In stark contrast is the codifferential $\delta$ which is built out of $d$ and the Hodge operator * and for which there is no such identity. However, some other identities involving $\delta$ can be found scattered in the literature, often in a wider setting. This short paper aims to gather such formulas in one compact space making it a useful toolbox for the practitioners of the field. Those formulas are extended further using a generalized interior product, producing new (up to our knowledge) formulas such as Eqs. (60)–(64). Our interest in such formulas came as a by-product of our studies on restrictions of differential operators from flat ambient space to cosmological spaces, more precisely computations needed for the generalization to $p$-form of earlier work[1].
Section II deals with the relationship between the codifferential $\delta$ and the Schouten-Nijenhuis bracket brought back to forms, facts known for a long time once cast in the framework of Gertenhaber algebras. Some explicit and little known formulas are gleaned along the way. In Sec. III the interior product $i_{v}$ is generalized in a somewhat natural manner, it provides a set of new identities of their own and in peculiar it involves naturally the Schouten-Nijenhuis bracket. A set of noteworthy formulas involving the differential operators $d$, $\mathscr{L}_{v}$, $\delta$ and $\square$ are recorded. Our conventions regarding differential geometry follow those of Ref. 2 and needed formulas are briefly recalled in appendix A. The outline of some longer calculations are gathered in appendix B in order not to clutter the main body of the paper.
@@ II The Schouten-Nijenhuis bracket on forms and its properties
Schouten[3, 4], and then Schouten-Nijenhuis[5, 6], bracket came from the natural endeavor in differential geometry of generalizing the Lie derivative not just along vector but also along tensors. In particular it is the unique extension generalizing the Lie derivative on multivectors[7, 8] (i.e. the sections of $\bigwedge^{p}TM$). In addition to this, it fulfills a wealth of algebraic relations making it one of the simplest incarnation of a Gerstenhaber algebra[9, 10, 11]. In this section we expose how, once brought back to forms, it arises out of the codifferential of the product of two forms. The bracket can be viewed from many angles[12, 13], in particular it plays a key role in Poisson geometry[14, 15]. For additional material regarding Schouten brackets and its generalizations we refer the reader to Ref. 16 for a beautiful overview.
@@
$\llbracket\lambda,\beta\rrbracket=\flat(\mathscr{L}_{\sharp\lambda}\sharp\beta),$ $\lambda\in\Omega^{1},\beta\in\Omega^{b}.$ (8)
@@
$\llbracket\lambda,\beta\rrbracket=(\mathscr{L}_{\sharp\lambda}\tilde{g})^{mn}j_{m}i_{n}\beta+\mathscr{L}_{\sharp\lambda}\beta,$ (9)
where $\beta$ is a $b$-form, $\tilde{g}$ is the corresponding metrics on forms and $j_{m}$, $i_{n}$ stand for $j_{e_{m}}$ and $i_{e_{n}}$. When $\sharp\lambda$ is a Killing vector it simplifies to $\llbracket\lambda,\beta\rrbracket=\mathscr{L}_{\sharp\lambda}\beta$.
@@
and makes the Gerstenhaber algebra an exact Gerstenhaber algebra, that is a Batalin-Vilkovisky algebra[19, 20]. Equation (29), or Eq. (19), has been known for a long time[21, 22] (see also Ref. 15) and can be recovered from Gerstenhaber original work[9] through Hochschild-Kostant-Rosenberg theorem[23], or be found by other means[24]. As the bracket is generated by $\delta$ it behaves nicely with respect to it, namely:
@@
$\llbracket\alpha,\beta\rrbracket=-(\nabla_{m}\alpha)\wedge i^{m}\beta+(-1)^{a+1}(i^{m}\alpha)\wedge\nabla_{m}\beta,$ (31)
@@ III A generalized interior product and its interplay with the Schouten-Nijenhuis bracket
for $\alpha\in\Omega^{a}$, $\beta\in\Omega^{b}$ and $\gamma\in\Omega^{c}$ forms, $\phi$ a function and $\omega$ the volume-form. For $\alpha,\beta\in\Omega^{p}$ two forms of the same degree one recovers the product between $p$-forms:
@@ Appendix A Overview of needed formulas of differential geometry
We recall that the interior product appears in the Lie derivative through Cartan formula: $\mathscr{L}_{v}=i_{v}d+di_{v}$.
@@
see say App. A of Ref. 1 for additional identities. The generalized interior product stems from Eq. (75).